OpenStudy (somegirl1999):

Find the height of a triangle with area = 110 ft2 and base = 10 ft. (1 point) 5.5 ft 22 ft 550 ft 11 ft

OpenStudy (whpalmer4):

\[A=\frac{1}2bh\] You know \(A\) and \(b\), so rearrange the equation to solve for \(h\)

OpenStudy (somegirl1999):

ok thanx

OpenStudy (whpalmer4):

As this is a multiple choice test, you coul also just try the answer choices in the formula and see which one fits.

OpenStudy (somegirl1999):

oh i see kk thanx

OpenStudy (somegirl1999):

it is multiple choice the answers are below the question lol

OpenStudy (whpalmer4):

I understand that. I'm helping you learn how to make the right choice. I don't need the answers to find the answer myself...

OpenStudy (somegirl1999):

i see lol srry i meant to close it

OpenStudy (whpalmer4):

Do you know which answer is correct?

OpenStudy (somegirl1999):

no not relly my dad is the math person in the family i was never good at it

OpenStudy (whpalmer4):

Do you want to see how to solve that one? There's nothing particularly difficult here.

OpenStudy (somegirl1999):

yes plz

OpenStudy (whpalmer4):

Great answer! :-) As I said earlier, the area of a triangle is given by \[A= \frac{1}2bh\] where \(A\) is area, \(b\) is the length of the base, and \(h\) is the height. We need to find \(h\) for this problem, so here are the basic steps to rearrange the formula to give \(h\) in terms of the other values: Multiply both sides of the equation by 2 \[2A = 2*\frac{1}2bh\]\[2A = bh\]Divide both sides by \(b\)\[\frac{2A}b = \frac{bh}{b}\]\[\frac{2A}b = h\]Switch sides\[h=\frac{2A}b\] Now we can plug in our known values: \[h=\frac{2*110}{10} =\]What do you get?

OpenStudy (somegirl1999):

i bet i got this wrong but i got 2200 tht is wrong isnt it srry like i said i am not good at this

OpenStudy (somegirl1999):

@whpalmer4

OpenStudy (whpalmer4):

2*110 = 220. 220/10 = 22... (sorry, I didn't notice that you had responded!)

OpenStudy (somegirl1999):

Lol its ok thank you !